\begin{thebibliography}{10}

\bibitem{Aba}
N.~Abatangelo and M.~Cozzi.
\newblock An elliptic boundary value problem with fractional nonlinearity.
\newblock {\em SIAM J. Math. Anal.}, 53(3):3577--3601, 2021.

\bibitem{Salo}
S.~Alarcón and A.~Quaas.
\newblock Large number of fast decay ground states to matukuma-type equations.
\newblock {\em J. Differential Equations}, 248:866--892, 2010.

\bibitem{MR1696454}
A.~Ambrosetti, J.~Garcia~Azorero, and I.~Peral.
\newblock Perturbation of {$\Delta u+u^{(N+2)/(N-2)}=0$}, the scalar curvature
  problem in {${\bf R}^N$}, and related topics.
\newblock {\em J. Funct. Anal.}, 165(1):117--149, 1999.

\bibitem{Mal}
A.~Ambrosetti and A.~Malchiodi.
\newblock {\em Perturbation Methods and Semilinear Elliptic Problems on
  $\mathbb{R}^n$}.
\newblock Progress in Mathematics. 2006.

\bibitem{Bahri}
A.~Bahri and J.~Coron.
\newblock On a nonlinear elliptic equation involving the critical sobolev
  exponent: The effect of the topology of the domain.
\newblock {\em Comm. Pure Appl. Math.}, 41:253--294, 1988.

\bibitem{Bamon}
R.~Bamón, I.~Flores, and M.~del Pino.
\newblock Ground states of semilinear elliptic equations: A geometric approach.
\newblock {\em Ann. Inst. H. Poincaré Anal. Non Linéaire}, 17:551--581, 2000.

\bibitem{Barles1}
G.~Barles, E.~Chasseigne, A.~Ciomaga, and C.~Imbert.
\newblock Lipschitz regularity of solutions for mixed integro-differential
  equations.
\newblock {\em J. Differential Equations}, 252:6012--6060, 2012.

\bibitem{Barles2}
G.~Barles, E.~Chasseigne, and C.~Imbert.
\newblock On the dirichlet problem for second order elliptic
  integro-differential equations.
\newblock {\em Indiana Univ. Math. J.}, 57(1):213--246, 2008.

\bibitem{bego}
B.~Barrios and A.~Quaas.
\newblock The sharp exponent in the study of the nonlocal hénon equation in
  $\mathbb{R}^n$: a liouville theorem and an existence result.
\newblock {\em Calc. Var. Partial Differential Equations}, 59(114), 2020.

\bibitem{Bernard}
C.~Bernard.
\newblock Regularity of solutions to the fractional laplace equation.
\newblock Preprint, 2014.

\bibitem{Bia1}
S.~Biagi, S.~Dipierro, E.~Valdinoci, and E.~Vecchi.
\newblock Semilinear elliptic equations involving mixed local and nonlocal
  operators.
\newblock {\em Proc. Roy. Soc. Edinburgh Sect. A}, 151(5):1611--1641, 2021.

\bibitem{Bia2}
S.~Biagi, S.~Dipierro, E.~Valdinoci, and E.~Vecchi.
\newblock Mixed local and nonlocal elliptic operators: regularity and maximum
  principles.
\newblock {\em Comm. Partial Differential Equations}, 47(3):585--629, 2022.

\bibitem{Bia}
S.~Biagi, S.~Dipierro, E.~Valdinoci, and E.~Vecchi.
\newblock A faber-krahn inequality for mixed local and nonlocal operators.
\newblock {\em J. Anal. Math.}, 150:405--448, 2023.

\bibitem{Biswas}
A.~Biswas, M.~Modasiya, and A.~Sen.
\newblock Boundary regularity of mixed local-nonlocal operators and its
  applications.
\newblock {\em Annali di Matematica Pura ed Applicata}, 202:679--710, 2023.

\bibitem{BonforteVa}
M.~Bonforte, P.~Ibarrondo, and M.~Ispizua.
\newblock The {C}auchy-{D}irichlet problem for singular nonlocal diffusions on
  bounded domains.
\newblock {\em Discrete Contin. Dyn. Syst.}, 43(3-4):1090--1142, 2023.

\bibitem{MR1412438}
K.~J. Brown and N.~Stavrakakis.
\newblock Global bifurcation results for a semilinear elliptic equation on all
  of {$\bold R^N$}.
\newblock {\em Duke Math. J.}, 85(1):77--94, 1996.

\bibitem{Caff1}
L.~A. Caffarelli, B.~Gidas, and J.~Spruck.
\newblock Asymptotic symmetry and local behavior of semilinear elliptic
  equations with critical {S}obolev growth.
\newblock {\em Comm. Pure Appl. Math.}, 42(3):271--297, 1989.

\bibitem{Campos}
J.~Campos.
\newblock “bubble-tower” phenomena in a semilinear elliptic equation with
  mixed sobolev growth.
\newblock {\em Nonlinear Anal.}, 68:1382--1397, 2008.

\bibitem{Chen}
W.~Chen and C.~Li.
\newblock Classification of solutions of some nonlinear elliptic equations.
\newblock {\em Duke Math. J.}, 3(3):615--622, 1991.

\bibitem{Davila}
J.~Davila, M.~del Pino, and M.~Musso.
\newblock The supercritical lane–emden–fowler equation in exterior domains.
\newblock {\em Commun. Partial Differential Equations}, 32(8):1225--1243, 2007.

\bibitem{MR3631323}
L.~M. Del~Pezzo and A.~Quaas.
\newblock A {H}opf's lemma and a strong minimum principle for the fractional
  {$p$}-{L}aplacian.
\newblock {\em J. Differential Equations}, 263(1):765--778, 2017.

\bibitem{manuel}
M.~del Pino, P.~Felmer, and M.~Musso.
\newblock Two-bubble solutions in the super-critical bahri-coron’s problem.
\newblock {\em Calc. Var. Partial Differential Equations}, 16(2):113--145,
  2003.

\bibitem{MR2944369}
E.~Di~Nezza, G.~Palatucci, and E.~Valdinoci.
\newblock Hitchhiker's guide to the fractional {S}obolev spaces.
\newblock {\em Bull. Sci. Math.}, 136(5):521--573, 2012.

\bibitem{Dipierro2024}
S.~Dipierro, X.~Su, E.~Valdinoci, and J.~Zhang.
\newblock Qualitative properties of positive solutions of a mixed order
  nonlinear {S}chr\"odinger equation.
\newblock {\em Discrete Contin. Dyn. Syst.}, 45(6):1948--2000, 2025.

\bibitem{Souplet}
L.~Dupaigne, B.~Sirakov, and P.~Souplet.
\newblock A liouville-type theorem for the lane–emden equation in a
  half-space.
\newblock {\em Int. Math. Res. Not.}, 2022(12):9024--9043, 2022.

\bibitem{MR4349926}
Y.~Fang, B.~Shang, and C.~Zhang.
\newblock Regularity theory for mixed local and nonlocal parabolic
  {$p$}-{L}aplace equations.
\newblock {\em J. Geom. Anal.}, 32(1):Paper No. 22, 33, 2022.

\bibitem{Felmer}
P.~Felmer, A.~Quaas, and M.~Tang.
\newblock On the complex structure of positive solutions to matukuma-type
  equations.
\newblock {\em Ann. Inst. H. Poincaré Anal. Non Linéaire}, 26:869--887, 2009.

\bibitem{MR2917408}
F.~Ferrari and I.~E. Verbitsky.
\newblock Radial fractional {L}aplace operators and {H}essian inequalities.
\newblock {\em J. Differential Equations}, 253(1):244--272, 2012.

\bibitem{fil}
C.~D. Filippis and G.~Mingione.
\newblock Gradient regularity in mixed local and nonlocal problems.
\newblock {\em Math. Ann.}, 388:261--328, 2024.

\bibitem{Floer}
A.~Floer and A.~Weinstein.
\newblock Nonspreading wave packets for the cubic schrödinger equation with a
  bounded potential.
\newblock {\em J. Funct. Anal.}, 69:397--408, 1986.

\bibitem{Ga}
P.~Garain and J.~Kinnunen.
\newblock On the regularity theory for mixed local and nonlocal quasilinear
  elliptic equations.
\newblock {\em Trans. Amer. Math. Soc.}, 375(8):5393--5423, 2022.

\bibitem{Gidas}
B.~Gidas and J.~Spruck.
\newblock Global and local behavior of positive solutions of nonlinear elliptic
  equations.
\newblock {\em Comm. Pure Appl. Math.}, 34(4):525--598, 1981.

\bibitem{Lieb}
E.~H. Lieb.
\newblock Sharp constants in the {H}ardy-{L}ittlewood-{S}obolev and related
  inequalities, 1983.

\bibitem{Ni}
C.~Lin and W.-M. Ni.
\newblock A counterexample to the nodal domain conjecture and a related
  semilinear equation.
\newblock {\em Proc. Amer. Math. Soc.}, 102:271--277, 1988.

\bibitem{MR232968}
W.~Magnus, F.~Oberhettinger, and R.~P. Soni.
\newblock {\em Formulas and theorems for the special functions of mathematical
  physics}, volume Band 52 of {\em Die Grundlehren der mathematischen
  Wissenschaften}.
\newblock Springer-Verlag New York, Inc., New York, enlarged edition, 1966.

\bibitem{varational1}
A.~Maione, D.~Mugnai, and E.~Vecchi.
\newblock Variational methods for nonpositive mixed local–nonlocal operators.
\newblock {\em Fract. Calc. Appl. Anal.}, 26:943--961, 2023.

\bibitem{MR1938656}
A.~M. Micheletti and A.~Pistoia.
\newblock Existence of blowing-up solutions for a slightly subcritical or a
  slightly supercritical non-linear elliptic equation on {$\Bbb R^n$}.
\newblock {\em Nonlinear Anal.}, 52(1):173--195, 2003.

\bibitem{MR1219814}
W.-M. Ni and I.~Takagi.
\newblock Locating the peaks of least-energy solutions to a semilinear
  {N}eumann problem.
\newblock {\em Duke Math. J.}, 70(2):247--281, 1993.

\bibitem{Poho}
S.~I. Poho\v{z}aev.
\newblock On the eigenfunctions of the equation {$\Delta u+\lambda f(u)=0$}.
\newblock {\em Dokl. Akad. Nauk SSSR}, 165:36--39, 1965.

\bibitem{ros}
X.~Ros-Oton and J.~Serra.
\newblock The pohozaev identity for the fractional laplacian.
\newblock {\em Arch. Ration. Mech. Anal.}, 213:587--628, 2014.

\bibitem{ros2}
X.~Ros-Oton and J.~Serra.
\newblock Nonexistence results for nonlocal equations with critical and
  supercritical nonlinearities.
\newblock {\em Comm. Partial Differential Equations}, 40:115--133, 2015.

\bibitem{val}
R.~Servadei and E.~Valdinoci.
\newblock Variational methods for non-local operators of elliptic type.
\newblock {\em Discrete Contin. Dyn. Syst.}, 33(5):2105--2137, 2013.

\bibitem{Stein}
E.~M. Stein.
\newblock {\em Singular integrals and differentiability properties of
  functions}.
\newblock Princeton Mathematical Series, No. 30. Princeton University Press,
  Princeton, NJ, 1970.

\bibitem{MR4492518}
X.~Su, E.~Valdinoci, Y.~Wei, and J.~Zhang.
\newblock Regularity results for solutions of mixed local and nonlocal elliptic
  equations.
\newblock {\em Math. Z.}, 302(3):1855--1878, 2022.

\bibitem{Talenti}
G.~Talenti.
\newblock Best constant in {S}obolev inequality.
\newblock {\em Ann. Mat. Pura Appl. (4)}, 110:353--372, 1976.

\bibitem{Yamabe}
H.~Yamabe.
\newblock On a deformation of {R}iemannian structures on compact manifolds.
\newblock {\em Osaka Math. J.}, 12:21--37, 1960.

\end{thebibliography}

\documentclass[11pt]{amsart}

	\usepackage{amsmath}
	\usepackage{amssymb}
	\usepackage{graphicx,color}
	\usepackage{amsxtra}
	\usepackage{mathtools}
	\usepackage{enumerate}
	\usepackage{nicefrac}
	\mathtoolsset{showonlyrefs} 
	\usepackage{ mathrsfs }
	\usepackage{tikz}
	\usetikzlibrary{positioning, calc}
	\usetikzlibrary{shadings}
	\usepackage{color}
	\usepackage{hyperref}
	\hypersetup{
	  colorlinks   = true, 
	  urlcolor     = blue, 
	  linkcolor    = blue, 
	  citecolor   = red 
	}

	\newtheorem{theorem}{Theorem}[section]
	\newtheorem{lemma}{Lemma}[section]

	\theoremstyle{definition}
	\newtheorem{definition}{Definition}[section]
	\newtheorem{remark}{Remark}[section]

	\DeclareMathOperator{\Ker}{Ker}
	\DeclareMathOperator{\spn}{span}
	\newcommand {\LL}{L_{a,b}}
	
    \title[Mixed local and nonlocal Lane-Emden equation]
    {Mixed local and nonlocal laplacian without standard critical exponent for 
    Lane-Emden equation}

   \author[B. Barrios]{Bego\~na Barrios}
	\address{Bego\~na Barrios \hfill\break\indent
		Departamento de An\'{a}lisis Matem\'{a}tico e IMAULL
		\hfill \break \indent Universidad de La Laguna
		\hfill \break \indent C/. Astrof\'{\i}sico Francisco S\'{a}nchez s/n, 
		38200 -- La Laguna, SPAIN}
	\email{bbarrios@ull.edu.es}
    
    \author[L. M. Del Pezzo]{Leandro M. Del Pezzo}
	\address{Leandro M. Del Pezzo \hfill\break\indent
            IESTA --Facultad de Ciencias Económicas y de Administración 	
            \hfill\break\indent
            Universidad de la República\hfill\break\indent
            Av. Gonzalo Ramírez 1926, 11200 Montevideo,\hfill\break\indent
            Departamento de Montevideo - Uruguay.}
	\email{leandro.delpezzo@fcea.edu.uy}

	\author[A. Quaas]{Alexander Quaas}
	\address{A. Quaas
       \hfill\break\indent Departamento de Matem\'atica, 
       Universidad T\'ecnica Federico Santa Mar\'ia 
       \hfill\break\indent Casilla V-110, Avda. Espa\~na, 1680 -- 
       Valpara\'iso, CHILE.}
    \email{alexander.quaas@usm.cl}

\begin{document}
    
   \begin{abstract}
 		In this paper, we investigate a mixed elliptic equation 
 		involving both local and nonlocal Laplacian operators, 
 		with a power-type nonlinearity. Specifically, we consider a 
 		Lane-Emden type equation of the form
        \[
            -\Delta u + (-\Delta)^s u 
				= u^p,\quad\mbox{ in }\mathbb{R}^n.
        \]
		where the operator combines the classical Laplacian and 
		the fractional Laplacian. We establish the existence 
		of solutions for exponents slightly below the critical local 
		Sobolev exponent, that is, for
		\(p < \frac{n+2}{n-2}\), with \(p\) close to
        \( \frac{n+2}{n-2}\).
        
		Our results show that, due to the interaction between the local 
		and nonlocal operators, this mixed Lane-Emden–Fowler equation 
		does not admit a critical exponent in the traditional sense. 
		The existence proof is carried out using a Lyapunov–Schmidt 
		type reduction method and, as far as we know, provide the first example of an elliptic operator
	for which the duality between critical exponents fails. 

    \end{abstract}
   \maketitle 
\section{Introduction}
	
	The objective of this paper is to explore critical exponents 
	associated with a Lane-Emden type equation 
	involving a mixed elliptic operator of the form
    \[
           \LL(u)\coloneqq -a\Delta u +b (-\Delta)^s u = u^p,
	\]
	with $a,b\in [0,\infty),$ with at least one of them strictly 
	positive. As is well known the fractional Laplacian is defined, up to a 
	normalization constant (which we omit for brevity), by a singular integral
	 \[
	 	(-\Delta)^s u(x)\coloneqq \lim_{r\to 0^+}
	 	\int_{\mathbb{R}^n\setminus B(x,r)}
		 \frac{u(x)-u(y)}{|x-y|^{n+2s}}dy.
	 \] 
We have to mention that the interest in problems involving mixed operators
has increased significantly in recent years. This is partly due to their ability
to capture phenomena where both local and nonlocal effects coexist that arises naturally in models combining classical
diffusion (governed by Brownian motion) and anomalous diffusion (described by
Lévy-type processes). Such a combination is relevant in diverse applications, including
materials with heterogeneous microstructure, financial models with both local volatility
and jumps. In biological systems where individuals may move via both short- and long-range interactions these type of operators appear too. From a theoretical perspective, these problems pose interesting challenges due to the competing regularity and scaling behaviors of the local and nonlocal parts.	Some non exhaustive list reference in the topic could be \cite{Barles1, Barles2,Aba, Bia, 
	Bia1,Bia2,Biswas,fil,Ga}. 
	
About the notion of critical exponents, it is well known that the critical exponent in the whole space associated to
    \[
 		F(u)= u^p,
    \]
    where \(F\) is a (possibly nonlinear) elliptic operator, 
    local or nonlocal in nature, is defined by
$$
		p_*(F)\coloneqq \inf \{p\in \mathbb{R}\colon
			\mbox{{\small there exists a positive 
				solution solution to} } F(u)=u^p \mbox{ {\Small in }} 		
					\mathbb{R}^n \}.
$$
It is also known by the community that Lane-Emden type equations at its critical exponent play a central 
	role in the study of PDE's problems as illustrated the classical case 
	$F=-\Delta$ which connects the Yamabe problem 
	and sharp constants in Sobolev inequalities. See for instance
	\cite{Yamabe,Talenti,Lieb}. 

On the other hand, the critical exponent in a bounded
	(smooth) domain is defined by
  	$$
		p^*(\Omega, F)\coloneqq
			\sup \{p\in \mathbb{R}\colon\mbox{ {\small there exists a 
			positive solution to} } F(u)= u^p 
			\mbox{ {\Small in} } \Omega \}.
	$$
 	
 	For the classical Laplacian $F=-\Delta$  and when $\Omega$  
 	is star-shaped, it is well known that a form of {\it duality}
 	appears, that is,
 	\[
 		p^*(\Omega, -\Delta)=p_*(-\Delta)=p_1\coloneqq\frac{n+2}{n-2},
 	\] 
	see \cite{Poho, Caff1, Chen, Gidas}.
	In the radial case,  
	this duality of the exponent (that is 
	$p^*(B_R, -\Delta)=p_*(-\Delta)$)  is clear, 
	since a solution $u_\gamma(r)$, $r \in (0,+\infty)$ is a 
	solution of an ODE with positive initial condition 
	$u(0)=\gamma>0$ 
	that cross or not the  $r$-axes. In other words,  if 
	$1<p<p_1$ then there exists $r_0$ such  $u_\gamma(r_0)=0,$ 
	and if $p\geq p_1$ then $u_\gamma(r)>0$ for all 
	$r \in (0,+\infty).$ By scaling $\gamma$ can 
	be an arbitrary value of $\mathbb{R}^+$. The Emden-Fowler transformation is a classical change of variable 
	to study this types of equations and gives the radial duality 
	described above.
 
	\medskip

	Thus, it seems natural to find some domains 
	$\Omega$ with non-trivial topology (not star-shaped) such that 
  	$p^*(\Omega, -\Delta)>p_1$. See for instance \cite{manuel} and also the starting 
	point in \cite{Bahri} at the critical exponent $p_1$.  The case of the exterior domain is discussed for example in 
	\cite{Davila} and the case of half space in \cite{Souplet}. 
	These types of domains 
	are interesting for mixed operators but 
	these cases fall outside the scope of this work in which we focus on the complete Euclidean space.

	\medskip
  
	Returning to star-shaped domains or the whole space, 
	similar results than the local ones, hold when considering the purely 
	nonlocal operator	
	$F=L_{0,1}=(-\Delta)^s$ (i.e. \(a=0\) and \(b=1\)). 
	In fact, if $\Omega$ is star-shaped then we also 
	have duality of the exponents and  
	\[
		p^*(\Omega, (-\Delta)^s)=p_*((-\Delta)^s)
		=p_s\coloneqq\frac{n+2s}{n-2s},
	\]
	(see \cite{ros,val,Chen}). Variational methods guarantee existence of solutions
	in any regular bounded domain $\Omega$ for 	
	the subcritical range $1<p<p_s,$ when $\Omega$ is
	not necessarily star-shaped, see \cite{val}. Other results posed in the pure nonlocal setting
regarding critical can be found in, for example,  
	\cite{bego}, where a H\'enon-type equation is discussed.
	
	{Some other equations that have sub and 
	supercritical phenomenas similar to the problem  
	\eqref{elproblema} are for instance,
	$$
		\Delta u+u^p+u^q=0\quad\text{in }\quad\mathbb{R}^n,\, 1<p<p_1<q,
	$$
	studied in the seminal work of \cite{Ni}. For that equation existence of positive 
	solution with fast decay are showed in  \cite{Bamon}. See also \cite{Campos}. }
	
	Another closely connected problem with 
	sub and supercritical 
	behavior is the Matukuna type problem of the form 
	$$
		\Delta u +K(r)u^p=0,\, \text{in }\mathbb{R}^n,
	$$
	studied in \cite{Felmer} where fast decay solution 
	exists for special $K$ 
	having sub and super critical behavior. 
	Moreover,  the case where $K$ is a mixed Henon type 
	with sub and supercritical behavior is studied in 
	\cite{Salo}. Fast decay solutions also exist.

	\medskip

	Returning to the case of mixed operators,  
	if $\Omega$ is a star-shaped domain and \(a,b>0\)
	it has been shown in Theorem 1.3 in \cite{ros2}( see also example iv 
	in Section 2) that $p^*(\Omega, \LL) =p_1$. This value of the critical exponent also 
	appears in a Brezis–Nirenberg type result \cite{Bia2}. Using 
	variational methods, existence results in bounded domains 
	have been established for the subcritical range
	$1<p<p_1$ in \cite{varational1} even when $b<0$.
	
	\medskip
	
	This leads to the  natural {\bf conjecture} for star-shaped domains
	\[
		p_*(\LL)=p^*(\Omega, \LL)=p_1,
	\] 
	that affirms that also the duality holds for 
	the mixed operator. 
	
	\medskip
	
	The main objective of the present work is to establish that the  
	{\bf conjecture} is false and
	that, in fact, if $\Omega$ is a star-shaped domain it holds
	\[
		p_*(\LL)<p^*(\Omega, \LL)=p_1,
	\]
	providing the first example of an elliptic operator $\LL$
	for which the duality between critical exponents fails.  
	Before detailing the concrete unexpected result we are able to prove,  
	we state a related conjecture that read as follows.

	\medskip
	\begin{center}
	{\bf Conjecture:} ${p_{*}(\LL)}=p_s$.
	\end{center}
	\medskip

	The proof of the previous conjecture need to be split into two parts. 
	
	\begin{itemize}
	\item[i)] If $1<p<p_s$ one may expect that there is no solution to 
	\begin{align}\label{ecu}
	\LL u=u^p\, \mbox{ in }\mathbb{R}^N.
	\end{align}
	We notice that if we assume some extra 
	integrability conditions on the solutions at its gradient, this can be proved by applying the Pohozaev identity given in Remark 3.1 
	of \cite{ros2}.
	  
	\item [ii)] The 
	existence of solutions to \eqref{ecu} for $p=p_s+\varepsilon$ could be proved following a similar (but not straightforward at all)
	perturbation-type arguments of the used in the present work. We think that a different kind of functional framework should be used and the {\it bubble} function associated to $(-\Delta)^s$ may play an important role in the bifurcation procedure.
	 
	\end{itemize}
	\medskip
	
	\medskip
	Such critical exponents are tightly related to fundamental threshold phenomena, where the existence of positive solutions typically changes abruptly. Moreover, understanding the value of them for general operators $F$, including nonlocal or mixed ones like $-\Delta + (-\Delta)^s$, not only extends the classical theory but also sheds light on new regimes of nonlinearity, where local and nonlocal effects interact in nontrivial ways.

	Coming back to our objective, as we mentioned, we want to establish the existence 
	of radial solutions to
	 \begin{equation}
		 \label{elproblema}
	 		\begin{cases}
	 			-\Delta u+ (-\Delta)^s u= n(n-2)u^{p_\varepsilon} 
					 &\text{in }\mathbb{R}^n,\, p_{\varepsilon}:=p_1-\varepsilon,\, n\geq 3,\\
					 u>0  &\text{in }\mathbb{R}^n
			 \end{cases}
		 \end{equation}
	where for simplicity,
	without loss of generality we have taken
	\[
		a=b=\frac{1}{n(n-2)}  
	\]
	in $L_{a,b}$, and  \(\varepsilon>0\) 
	will be a small parameter. It is clear that  \(u_\varepsilon\) is a solution of 
	 \eqref{elproblema} if and only if
	 \[
	 u_{\delta, \varepsilon}(x)\coloneqq \delta^{\frac2{p_\varepsilon-1}} 
	 u_\varepsilon(\delta x),
	 \text{ with }\delta>0,
	 \]
	 is a solution of
	 \begin{equation}
	 \label{elproblemaDelta}
	 \begin{cases}
	 -\Delta u+ \delta^{2(1-s)}(-\Delta)^s u= 
	 n(n-2)u^{p_\varepsilon} &\text{in }\mathbb{R}^n,\\
	 u>0  &\text{in }\mathbb{R}^n,
	 \end{cases}
	 \end{equation}
	 in a suitable energy space (see \eqref{elespacio} below). Therefore, since it is well known (see \cite{Talenti}) that the local problem 
	 \begin{equation}
	 \label{elproblemalocalconp0}
	 \begin{cases}
		 -\Delta u= n(n-2)u^{p_1} &\text{in }\mathbb{R}^n,\\
		 u>0  &\text{in }\mathbb{R}^n, 
	 \end{cases}
	 \end{equation}
	 has the family of Talenti functions
	 (scale by the initial condition)
\begin{align}\label{bubble}
	 	U_{\delta} (x)\coloneqq\delta^{\frac{n-2}2} 
	 	U\left(\frac{x}{\delta}\right),\, U(x)\coloneqq \frac1{(1+|x|^2)^{\frac{n-2}2}},\, \delta>0, y\in\mathbb{R}^n,
\end{align}
as unique solutions, one of our goals is to show the following.
	
	\begin{theorem}	\label{ElTeorema}
	 	 Let \(n\geq 3\) and $\varepsilon>0$. There exists $\delta(\varepsilon)$ such that the problem
	 	 \eqref{elproblemaDelta} has a radial solution of the form
		 \[
			 z_\varepsilon(x)= U(x) +\phi_\varepsilon(x),
		 \]
		where $\phi_\varepsilon \to 0$ in
		$\mathcal{D}^{1,2}(\mathbb{R}^n)$ and $U$ is given in \eqref{bubble}.
	 	In addition, by rescaling back, there exists a solution 
	 	of \eqref{elproblema} that concentrates at the origin.
	 \end{theorem}	
To prove it, we use a
	Lyapunov–Schmidt reduction method that has first used in \cite{Floer} in the 
	context of partial differential equations and has become a very 
	active area in PDEs since then. 
	For a review of some other perturbation methods 
	in $\mathbb{R}^n$ see the book \cite{Mal} and also 
	\cite{Campos,Salo}. In our particular case we follow the 
	Lyapunov–Schmidt reduction in the spirit of \cite{MR1938656} 
	wherethe difficulties that the nonlocal operator 
	introduces appear. The main idea of this approach is to solve 
	the problem in the orthogonal of the 
	kernel of the linearized operator denoted by $K$ with some 
	parameters (finite reduction) 
	and then find parameters so that there is 
	also a solution to the full problem. 
	The parameters are chosen so 
	the problem is always in the orthogonal space of the Kernel  
	$K$  so is a solution to the problem in all space (not only in the 
	orthogonal of the $K$).

			\medskip 
		The paper is organized as follows. In Section 2 we introduce the framework and the notation we will use along the work. Moreover we prove several auxiliar lemmas that will be needed to establish the proof of Theorem \ref{ElTeorema} that is detailed in Section 3. We want to highlight the Lemma \ref{taqui}, a key point in our approach, 
	which establishes that the
	$H^s(\mathbb{R}^n)$ semi-inner product of the bubble 
	solution with 
	the unique element of the Kernel $K$ 
	has a negative sign. From here we can find parameters to guarantee
	that the projected solution is a full 
	solution.

		
    \section{Problem Setting and Notations}\label{section.psn}
		
		Given \(s\in(0,1]\), \(\mathcal{D}^{s,2}(\mathbb{R}^n)\) 
		denotes the clausure of 
		\(C_0^\infty(\mathbb{R}^n)\) with respect to the norm
		\[
			\|u\|_{s,2}^2\coloneqq
			\begin{cases} 
			\displaystyle\int_{\mathbb{R}^n}\int_{\mathbb{R}^n} 
			\frac{|u(x)-u(y)|^2}{|x-y|^{n+2s}} dy\, dx{=[u]_{H^s}}&\text{if } s \in(0,1),\\[7pt]
			\displaystyle\int_{\mathbb{R}^n} |\nabla u|^2 \, \, dx{=[u]_{H^1}}  &\text{if } s=1.
			\end{cases}
		\]
		
		It is important to note that when it is equipped with the inner product
		\[
			\langle u,v\rangle_{s}\coloneqq
			\begin{cases}
				\displaystyle \int_{\mathbb{R}^n}\int_{\mathbb{R}^n} 
				\frac{(u(x)-u(y))(v(x)-v(y))}{|x-y|^{n+2s}} dy\, dx &\text{if } 
					s\in(0,1),\\[7pt]
			 	\displaystyle \int_{\mathbb{R}^n} \nabla u\nabla v \, dx  
			 	&\text{if } s=1,
			\end{cases}
		\]
		the space \(\mathcal{D}^{s,2}(\mathbb{R}^n)\) 
		becomes a Hilbert space. In the
		case \(s=1,\) for notational convenience we write 
		\(\langle \cdot,\cdot\rangle\) in place of 
		\(\langle \cdot,\cdot\rangle_1.\) We set 
\begin{align}\label{elespacio}
\mathcal{X}\coloneqq 
		\mathcal{D}^{1,2}(\mathbb{R}^n)\cap L^{\frac{2n}{n+2}}(\mathbb{R}^n),
\end{align}
		with the norm  
\begin{align}\label{lanorma}
			\|u\|\coloneqq \max
			\left\{\|u\|_{1,2},\|u\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}
			\right\},
\end{align}

		By the Nirenberg-Sobolev inequality and interpolation, we have the inclusion \(\mathcal{X}\subset L^q(\mathbb{R}^n)\) with
		\(q\in[2n/(n+2),2n/(n-2)].\) In particular, since \(2\in [2n/(n+2),2n/(n-2)],\) it follows that \(\mathcal{X}\subset L^2(\mathbb{R}^n).\)
		Consequently, by \cite[Proposition 2.1]{MR2944369},  we deduce
		\begin{equation}\label{relacion}
		    \mathcal{X}\subset \mathcal{D}^{s,2}(\mathbb{R}^n),
		\end{equation}
		for any \(s\in(0,1)\).
		
		\begin{definition}
			Given \(s\in(0,1)\), and \(\delta\ge 0,\) we define the operator
			\[
				\mathcal{I}_\delta\colon 
				L^{\frac{2n}{n+2}}(\mathbb{R}^n)\to\mathcal{D}^{1,2}(\mathbb{R}^n),
			\] 
			by setting
			\[
				\mathcal{I}_\delta(h)=u,
			\]
			if and only if \(u\) is a weak solution of
			\begin{equation}
				\label{eqconh}
				(-\Delta) u+ \delta^{2(1-s)}(-\Delta)^s u + \delta^{2(1-s)} u=h\quad\mbox{in}\quad \mathbb{R}^n.
			\end{equation}
			That is, \(u\in\mathcal{D}^{1,2}(\mathbb{R}^n)\) satisfies
			\[
				\langle u, v\rangle+\delta^{2(1-s)}
				\langle u, v\rangle_s +\delta^{2(1-s)} \int_{\mathbb{R}^n} uv \, dx
				=\int_{\mathbb{R}^n} hv \, dx,
			\]
			for any \(v\in\mathcal{D}^{1,2}(\mathbb{R}^n).\)
		\end{definition}
		
		By \cite[Lemma 6.9 and Corollary 6.7]{Bernard}, \cite[Section 3.3 in Chapter 5]{Stein} and \cite[Lemma 3.3]{Dipierro2024},
		we have the following result.
		
		\begin{lemma}\label{lem1}
			Let \(\delta\ge0\) and \(h\in L^{\frac{2n}{n+2}}(\mathbb{R}^n).\)
			If $u\in \mathcal{D}^{1,2}(\mathbb{R}^N)$ is a weak solution of \eqref{eqconh} 
			then \(u\in W^{2,\frac{2n}{n+2}}(\mathbb{R}^n)\) and there exists \(C=C(n,\delta)>0\) such that
			$$
				\|u\|_{W^{2,\frac{2n}{n+2}}(\mathbb{R}^n)}\leq  C \|h\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^{n})}. 
			$$	
		\end{lemma}

		\begin{remark}\label{remark:Ideltacont}
			As a consequence of Lemma~\ref{lem1}, the operator \(\mathcal{I}_\delta\)
			maps into the space \(\mathcal{X},\) that is
			\[
			\mathcal{I}_\delta\colon 
			L^{\frac{2n}{n+2}}(\mathbb{R}^n)\to\mathcal{X}.
			\]
			Moreover  \(\mathcal{I}_\delta\) is continuous. 
			Specifically, there exists a constant  \(C=C(n,\delta)\)
			such that
			\[
				\|\mathcal{I}_\delta(h)
				\|\le C\|h\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)},
			\]
			for all \(h\in L^{\frac{2n}{n+2}}(\mathbb{R}^n).\)
		\end{remark}
		
		Clearly \(u\) is a weak solution of \eqref{elproblemaDelta} if 
		and only if
		\begin{equation}\label{elproblemaDelta2}
				u=\mathcal{I}_\delta(g_{\varepsilon,\delta}(u)) 
				\text{ and } u\in\mathcal{X}.
		\end{equation}
		where 
\begin{align}\label{la_g}
			g_{\varepsilon,\delta}(t):= f_\varepsilon(t)  + \delta^{2(1-s)}t,
			\text{ with } 
			f_\varepsilon(t)=n(n-2) (t_+)^{p_\varepsilon},
\end{align}
		here \(t_+\) denotes the positive part of \(t.\) We also notice that
		
		\begin{lemma}
			If \(u\in\mathcal{X}\) is a 
			non-trivial weak solution of \eqref{elproblemaDelta2} (i.e. of \eqref{elproblemaDelta})
			then \(u>0\) a.e. in \(\mathbb{R}^n.\) 
		\end{lemma}
		\begin{proof}
Let  \(u\in\mathcal{X}\) be a non-trivial solution of \eqref{elproblemaDelta2}. Testing \eqref{elproblemaDelta2} with \(u_{-}(x)\coloneqq \max\{-u(x),0\}\geq 0,\) ($u=u_{+}-u_{-}$, $u_{+}=\max\{u,\, 0\}$ ), we clearly get
\begin{align}
\begin{split}\label{weak}
0&=-\|\nabla u_-\|_{1,2}^2+\delta^{2(1-s)}\langle u_{+}-u_{-},u_{-}\rangle_{s}\\
&=-\|\nabla u_-\|_{1,2}^2+\delta^{2(1-s)}\left(\langle u_{+},u_{-}\rangle_{s} -[u_{-}]^{2}_{s,2}\right),
\end{split}
\end{align}
so that $\langle u_{+},u_{-}\rangle_{s}\geq 0$ with in fact implies that $\langle u_{+},u_{-}\rangle_{s}=0$. Thus, from \eqref{weak} we obtain that $u_{-}\equiv C,\, C\in \mathbb{R}$ that, since $u\in\mathcal{X}$, implies $C=0$ as wanted. To get the strictly positivity of the solution, since we are working in weak sense, using a Logarithmic Lemma (see \cite{MR4349926}) and following the ideas developed in \cite[Lemma 3.3]{MR3631323} we conclude.

\end{proof}

		Our next goal is to derive estimates depending on the parameters \(\varepsilon\) and \(\delta.\)
		For this purpose, we need the following  \(L^q-\)estimate for the fractional Laplacian.
		
		\begin{lemma}
			\label{lemma:Lpesitmate}
			For any \(u\in W^{2,q}(\mathbb{R}^n),\) it holds that \((-\Delta)^su\in L^q(\mathbb{R}^n)\) and
			\[
				\|(-\Delta)^su\|_{L^q(\mathbb{R}^n)}\le C\|u\|_{W^{2,q}(\mathbb{R}^n)},
			\]
			for some positive constant \(C\)  depending only on \(n,s\) and \(q.\)   
		\end{lemma}
		\begin{proof}
			The proof follows ideas from \cite[Subsection 4.1.1]{MR4492518}.
			
			Let \(u\in W^{2,q}(\mathbb{R}^n).\) By the Minkowski inequality, we estimate
			\begin{align*}
				&\left(\int_{\mathbb{R}^n}
					\left|
						\int_{|y|<1}\frac{u(x+y)+u(x-y)-2u(x)}{|y|^{n+2s}} dy
					\right|^q\, dx\right)^{\frac1q}
				\le\\ 
				&\le \left(\int_{\mathbb{R}^n}\left(
					\int_{|y|<1}\int_0^1\int_0^\theta |D^2u(x+\eta y)| 
					d\eta d\theta \frac{dy}{|y|^{n+2(s-1)}}\right)^q\, dx\right)^{\frac1q}\\
				&\le \int_{|y|<1} \left(\int_{\mathbb{R}^n}
				\int_0^1\int_0^\theta |D^2u(x+\eta y)|^q d\eta d\theta \, dx\right)^{\frac1q}\frac{dy}{|y|^{n+2(s-1)}}\\
				&\le C(n,s)\|u\|_{W^{2,q}(\mathbb{R}^n)}.
			\end{align*}	
				
			For the remainder of the proof, that is for the integral estimate when $|y|\geq 1$ 
			we refer to the estimate of  \(III\) in the proof of Lemma 4.2 in \cite{MR4492518}.
			This completes the proof.
		\end{proof}
		
		\begin{lemma}\label{lemma:ordenconv}
		    Let \(s\in(0,1)\),  \(0<\varepsilon,\delta<1.\) Let
		    \(\phi\in \mathcal{X},\) and define
		    \[
		        w_{\varepsilon,\delta}
		        =\mathcal{I}_\delta(g_{\varepsilon,\delta}(U+\phi)),
		    \]
		    where $g_{\varepsilon,\delta}$ is given in \eqref{la_g} and $U$ is defined in \eqref{bubble}
		    Then, for \(\varepsilon,\) and \(\delta\) small enough, there exists a constant \(C>0\) independent 
		    of \(\varepsilon,\) and \(\delta\) such that
		    \[
		    	\|w_{\varepsilon,\delta}-U\|\le C\left(\|\phi\|^{{p_1-\varepsilon}}+\|\phi\|+\varepsilon+\delta^{2(1-s)}\right),
		    \]
		    for any \(\phi\in\mathcal{X}\) where $\|\cdot\|$ was given in \eqref{lanorma}.
		\end{lemma}
		\begin{proof}
			Let \(z\coloneqq w_{\varepsilon,\delta}-U.\) Then $z$ satisfies the equation
			\[
				-\Delta z+\delta^{2(1-s)}(-\Delta)^s z+
				 \delta^{2(1-s)}z=f_\varepsilon(U+\phi)-
				f_0(U)+ \delta^{2(1-s)}\phi- \delta^{2(1-s)}(-\Delta)^sU.
			\]
			Firstly we estimate each terms of the right hand side 
			\[
				H\coloneqq
				f_\varepsilon(U+\phi)-
				f_0(U)+ \delta^{2(1-s)}\phi- \delta^{2(1-s)}(-\Delta)^sU.
			\]
in the space \(L^{\frac{2n}{n+2}}(\mathbb{R}^n).\) For that we observe that by \cite[Remarks 2.21 and 2.22]{MR1938656}, 
			for sufficiently small \(\varepsilon,\) and \(\delta\)
			there exists a positive constant \(C\), independent 
			of \(\varepsilon,\) and \(\delta\), such that
			\begin{align*}
				\|f_\varepsilon(U+\phi)-&f_0(U)\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}=
				\|f_\varepsilon(U+\phi)-f_\varepsilon(U)-
				f^\prime_\varepsilon(U)\phi\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}\\
				&+\|f_\varepsilon(U)-f_0(U)\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}+
				\|f^\prime_\varepsilon(U)\phi\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}\\
				&\le C\left(\|\phi\|^{p_1-\varepsilon}+\varepsilon+\|\phi\|\right).
			\end{align*}
Moreover, since from \cite[Page 259]{MR2917408}, 
			we know that 
\begin{align}\label{Condicion1}
(-\Delta)^s U \in L^{\frac{2n}{n+2}}(\mathbb{R}^n),
\end{align}
we clearly get that		
\begin{align}\label{lio1}
\|H\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}\le 
				C\left({\|\phi\|^{p_1-\varepsilon}}+\|\phi\|+\varepsilon+\delta^{2(1-s)}\right),
\end{align}
			for some positive constant \(C,\) independent of \(\varepsilon\) and \(\delta,\) provided both are sufficiently small. Then, by Lemma \ref{lem1} we get that \(z\in W^{2,\frac{2n}{n+2}}(\mathbb{R}^n)\) and
			\[
				\|z\|_{L^{\frac{2n}{n+2}(\mathbb{R}^n)}}\le	\|z\|_{W^{2,\frac{2n}{n+2}}(\mathbb{R}^n)}
					\le C\left({\|\phi\|^{p_1-\varepsilon}}+\|\phi\|+\varepsilon+\delta^{2(1-s)}\right),
			\]
			where \(C\) is a positive constant independent of \(\varepsilon\) and \(\delta,\) provided both are sufficiently small.
			
			On the other hand, by H\"older and Nirenberg-Sobolev inequalities, by \eqref{lio1} we also obtain
			\begin{align*}
				\|z\|_{1,2}^2&+\delta^{2(1-s)}\langle z,z\rangle_s
				+ \delta^{2(1-s)}\|z\|_{L^2(\mathbb{R}^n)}^2
				= \int_{\mathbb{R}^n}
				Hz\, dx\\
				&\le \|H\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}\|z\|_{L^{\frac{2n}{n-2}}(\mathbb{R}^n)}\\
				&\le C\left({\|\phi\|^{p_1-\varepsilon}}+\|\phi\|+\varepsilon+\delta^{2(1-s)}\right)	\|z\|_{1,2}.
			\end{align*}
			Therefore
			\[
				\|z\|_{1,2}\le C\left({\|\phi\|^{p_1-\varepsilon}}+\|\phi\|+\varepsilon+\delta^{2(1-s)}\right),
			\]
			which concludes the proof. 		
		\end{proof}
We introduce the following fundamental definition.
		
		\begin{definition}
			Let \(\delta\ge0\). 
			The operator
			\(L_\delta\colon \mathcal{X}\to\mathcal{X}\) is given by
\begin{align}\label{Ele_delta}
				L_\delta(u)\coloneqq u- \mathcal{I}_\delta(g_{0,\delta}^\prime(U)u).
\end{align}
		\end{definition}
		
		 By \cite[Lemma 2.8]{MR1412438} we notice that \(L_0\) is a self-adjoint operator 
		and is a compact perturbation of the identity on \(\mathcal{X}\). We show now that the family of operators  \(L_\delta\) 
	    converges to \(L_0\) as \(\delta\to 0.\)

		\begin{lemma}\label{lema:limite}
			For every $u\in\mathcal{X}$ there exists $C>0$, independent of \(\delta\), such that 
				\[
					\|L_\delta(u)-L_0(u)\|\leq C \delta^{2(1-s)},
				\]
			provided as \(\delta\) is small enough.
		\end{lemma}
		\begin{proof}
			Let \(u\in\mathcal{X},\) and 
			\(w_\delta\coloneq L_\delta(u),\, \delta\geq 0.\)
			Then, by definition
\begin{align}\label{ec_aux}
-\Delta (u-w_0)={n(n+2)U^{\frac{4}{n-2}}}u.
\end{align}
			Since {\(0<U\in L^\infty(\mathbb{R}^n),\)} and \(u\in L^{\frac{2n}{n+2}}(\mathbb{R}^n)\),
			it follows from Lemma \ref{lem1} that \(u-w_0\in W^{2,\frac{2n}{n+2}}(\mathbb{R}^n).\) Moreover, 
			by Lemma \ref{lemma:Lpesitmate}, we also have that \((-\Delta)^s(u-w_0)\in L^{\frac{2n}{n+2}}(\mathbb{R}^n).\) 
			
			Let now \(z_\delta\coloneqq w_0-w_\delta\)  and \(\widetilde{H}\coloneqq (-\Delta)^s(w_0-u)+w_0 \in  L^{\frac{2n}{n+2}}(\mathbb{R}^n).\) 
			Since $w_\delta=u-v$ with
			\begin{align*}
			(-\Delta)v+\delta^{2(1-s)}(-\Delta)^s v +\delta^{2(1-s)} v=n(n+2)U^{\frac{4}{n-2}}u+\delta^{2(1-s)u},
			\end{align*}
			by \eqref{ec_aux} we get that \(z_\delta\) satisfies the equation
			\begin{equation}
				\label{eq.z}
					-\Delta z_\delta+\delta^{2(1-s)}(-\Delta)^sz_\delta+\delta^{2(1-s)}z_\delta= \delta^{2(1-s)}
					\widetilde{H}.
			\end{equation}
			By Lemma \ref{lem1}, we deduce that \(z_\delta\in W^{2,\frac{2n}{n+2}}(\mathbb{R}^n)\) and
			\[
					\|z_\delta\|_{ L^{\frac{2n}{n+2}}(\mathbb{R}^n)}\le\|z_\delta\|_{ W^{2,\frac{2n}{n+2}}(\mathbb{R}^n)}\le C\delta^{2(1-s)},
			\]
			where \(C>0\) is independent of \(\delta.\)
			
			Testing \eqref{eq.z} against \(z_\delta\) and using Nirenberg-Sobolev inequality, we get
			\[
				\|z_\delta\|_{1,2}^2\le\delta^{2(1-s)}\|\widetilde{H}\|_{L^{\frac{2n}{n+2}}(\mathbb{R})}
				\|z_\delta\|_{L^{\frac{2n}{n-2}}(\mathbb{R}^n)}
				\le C\delta^{2(1-s)}	\|z_\delta\|_{1,2},
			\]
			where \(C\) is a positive constant independent of \(\delta.\) This completes the proof.
		\end{proof}
		
		From now on, we restrict our analysis to the subspace 
\begin{align*}
			\mathcal{R}\coloneqq \{u\in\mathcal{X}\colon u \text{ is radial}\}.
\end{align*}
of radial functions. We have the following result whose proof we outline for completeness. 
		\begin{lemma}\label{lemaeq:ker}
			Let \(\delta\ge 0.\) If \(u\in\mathcal{R}\) then 
			\(L_\delta(u)\in \mathcal{R},\)
			that is \(L_\delta\bigr|_\mathcal{R}\colon\mathcal{R}
			\to \mathcal{R}.\)
			Moroever, 
			\begin{equation}\label{eq:ker}
	            \Ker L_0\bigr|_\mathcal{R}=\spn\{\psi\}
	        \end{equation}
			where
\begin{align}\label{la_psi}
				\psi(x)\coloneq x\cdot\nabla U+\frac{n-2}2 U= 
				\frac{n-2}2\frac{1-|x|^2}{(1+|x|^2)^{n/2}}.
\end{align}
		\end{lemma}
		\begin{proof}
		    The proof of
		    \(L_\delta\bigr|_\mathcal{R}\colon\mathcal{R}
			\to \mathcal{R},\) is given in \cite{Dipierro2024}. 
		    To show that \eqref{eq:ker} holds, 
		    see \cite[Lemma 3.12]{MR1696454}
		    or \cite[Lemma 4.2]{MR1219814}.
		\end{proof}
		\begin{remark}\label{remark.psi}
			If \(n\geq 3\) then $\psi\in {H^1(\mathbb{R}^{n})}$ 
			and therefore $\psi\in {H^s(\mathbb{R}^{n})}$.
		\end{remark}
		We define now
		\[
		    K\coloneqq\left\{v\in\mathcal{R}\colon 
		            \langle v,\psi\rangle_{1,2}=0
				\right\},
		\]
		and the continuous projection
		\[
		    \Pi\colon \mathcal{R}\to K,
		\]
		Moreover, for every \(\delta\ge 0\)  we define the operator  
		\(\widetilde{L}_\delta\colon K\to K\)
		as  
		\[
		    \widetilde{L}_\delta(u)
		    \coloneqq \Pi(L_\delta\bigr|_\mathcal{R}(u)).    
		\] 
		where $L_\delta$ was given in \eqref{Ele_delta}. According to  \cite{MR1938656}, 
		   the operator \(\widetilde{L}_0\) is invertible and 
		   \(\widetilde{L}_0^{-1}\) is continuous. Moreover we can shows that \(\widetilde{L}_\delta\) is invertible and 
		    \(\widetilde{L}_\delta^{-1}\) is continuous, as the following result reads.
\begin{lemma}\label{orz}
There exists \(\delta_0>0\) such that \(\widetilde{L}_\delta\) 
		    is invertible and \(\widetilde{L}_\delta^{-1}\) is continuous, for every
		    \(\delta\in(0,\delta_0)\) .
		\end{lemma}
		\begin{proof}
		   Using Lemma \ref{lema:limite}, is clear that $\widetilde{L}_\delta(u)=\widetilde{L}_0 (u)+
		      R_\delta(u)$ where 
		   \begin{equation}\label{eq:resto}
                \|R_\delta\|\le C\delta^{2(1-s)},
           \end{equation}
		  with \(C\)  being a positive constant independent of  \(\delta.\)
		   Then
		    \[
		      \widetilde{L}_\delta
		      = \widetilde{L}_0(Id+
		        \widetilde{L}_0^{-1} R_\delta ).
		   \]
		   Thus, by \eqref{eq:resto}, for sufficiently small
		   \(\delta,\) the operator
		   \(\widetilde{L}_\delta\) is invertible and
		   \[
		       \widetilde{L}_\delta^{-1}
		      = (Id+\widetilde{L}_0^{-1} R_\delta)^{-1}
		      \widetilde{L}_0^{-1}.
		   \]    
		\end{proof}

 We conclude this section with the following lemma that will be essential for proving our main result.
		
		\begin{lemma} \label{taqui}
		Let \(n>2\) and \(s\in(0,1)\). Then
		    \begin{equation}\label{Pistoia}
		        \int_{\mathbb{R}^n} \log(U) U^{p_1} \psi \, dx>0,
		    \end{equation}
		    and
		    \begin{equation}\label{chile241106}
		        \langle U,\psi \rangle_s<0,
		    \end{equation}
		    where $\psi$ was given in \eqref{la_psi}.
		\end{lemma}
        \begin{proof}
            To prove \eqref{Pistoia}, see the proof of Lemma 2.19 in \cite{MR1938656}.

            For \eqref{chile241106}, we use \cite[(6.5)]{MR2917408}, where it is 
            shown that
            \[
                (-\Delta)^s U(x) = C(n, s) F\left(\frac{n+2s}{2}, 
                \frac{n-2(1-s)}{2}, \frac{n}{2}, -|x|^2\right),
            \]
            where \( C(n, s) \) is a positive constant and \( F \) 
            denotes the Gaussian hypergeometric function. Then,
            \begin{equation}
                \label{eq:signo1}
                \begin{aligned}
                   &\langle U, \psi \rangle_s = C(n, s) \frac{n-2}{2} 
                    \int_{\mathbb{R}^n}
                    F\left(\frac{n+2s}{2}, \frac{n-2(1-s)}{2}, 
                    \frac{n}{2}, -|x|^2\right) \psi(x) \, \, dx \\
                    &= K(n, s) \int_0^{\infty} F\left(\frac{n+2s}{2}, 
                    \frac{n-2(1-s)}{2}, \frac{n}{2}, -\rho^2\right) 
                    \frac{1 - \rho^2}{(1 + \rho^2)^{n/2}} \rho^{n-1} \, d\rho \\
                    &= K(n, s) (I_1 + I_2),
                 \end{aligned}
            \end{equation}
            where \( K(n, s) \) is a positive constant, and we define
            \[
                I_1 \coloneqq \int_0^1 F\left(\frac{n+2s}{2}, \frac{n-2(1-s)}{2},
                \frac{n}{2}, -\rho^2\right) \frac{1 - \rho^2}
                {(1 + \rho^2)^{n/2}} \rho^{n-1} \, d\rho,
            \]
            and
            \[
                I_2 \coloneqq \int_1^{\infty} F\left(\frac{n+2s}{2}, 
                \frac{n-2(1-s)}{2}, \frac{n}{2}, -\rho^2\right) 
                \frac{1 - \rho^2}{(1 + \rho^2)^{n/2}} \rho^{n-1} \, d\rho.
            \]

            By making the substitution \( \rho = \tau^{-1} \), we have
            \begin{align*}
                I_2 &= \int_0^1 F\left(\frac{n+2s}{2}, \frac{n-2(1-s)}{2}, 
                \frac{n}{2}, -\tau^{-2}\right) 
                \frac{1 - \tau^{-2}}{(1 + \tau^{-2})^{n/2}} 
                \tau^{1-n} \frac{d\tau}{\tau^2} \\
                & = -\int_0^1 F\left(\frac{n+2s}{2}, \frac{n-2(1-s)}{2}, 
                \frac{n}{2}, -\tau^{-2}\right) \frac{1 - \tau^2}{(1 + 
                \tau^2)^{n/2}} \tau^{-3} \, d\tau.
            \end{align*}
            Then,
            \[
                I_1 + I_2 = \int_0^1 H(\rho) 
                \frac{1 - \rho^2}{(1 + \rho^2)^{n/2}} \, d\rho,
            \]
            where
            \begin{align*}
                H(\rho) = F&\left(\frac{n+2s}{2}, 
                \frac{n-2(1-s)}{2}, \frac{n}{2}, -\rho^2\right) \rho^{n-1} \\
                &- F\left(\frac{n+2s}{2}, \frac{n-2(1-s)}{2}, \frac{n}{2}, 
                - \rho^{-2}\right) \rho^{-3}.
            \end{align*}

            If
            \begin{equation}\label{eq:signoH}
                H(\rho) < 0 \quad \text{ in } (0, 1),
            \end{equation}
            then
            \[
                I_1 + I_2 < 0.
            \] 
            Thus, from \eqref{eq:signo1}, we conclude that \eqref{chile241106} 
            holds. To complete the proof, it remains to verify that 
            \eqref{eq:signoH} holds.

            By \cite[Section 2.5, page 54]{MR232968}, we have
            \begin{align*}
                &\frac{\Gamma\left(\frac{n-2(1-s)}{2}\right) 
                \Gamma(1-s)}{\Gamma\left(\frac{n}{2}\right)} H(\rho)=\\ 
                &=\int_0^1 
                \frac{t^{(n-2(2-s))/2}}{(1 - t)^s} 
                \left[ (1 + \rho^2 t)^{-(n+2s)/2} \rho^{n-1} 
                - (\rho^2 + t)^{-(n+2s)/2} \rho^{n+2s-3} \right] \, dt\\
                &=
                \int_0^1 \frac{t^{(n-2(2-s))/2}}{(1 - t)^s} 
                \frac{\rho^{n+2s-3}}{(\rho^2 + t)^{(n+2s)/2}} 
                \left[ \left( \frac{\rho^2 + t}{1 + \rho^2 t} \right)^{(n+2s)/2} 
                \rho^{2(1-s)} - 1 \right] 
                \, dt.
            \end{align*}
            Since
            \[
                \left( \frac{\rho^2 + t}{1 + \rho^2 t} \right)^{(n+2s)/2} 
                \rho^{2(1-s)} - 1 < 0 \quad \forall t, \rho \in (0, 1),
            \]
            we conclude that
            \[
                H(\rho) < 0 \quad \forall \rho \in (0, 1).
            \]        
        \end{proof}  
        
    \section{Existence}
    	In this section, we demonstrate the existence of a radial 
    	function \(\phi_{\varepsilon,\delta}\) such that 
    	\(U + \phi_{\varepsilon,\delta}\) satisfies equation 
    	\eqref{elproblemaDelta2}.
    	
    	\medskip
    	
    	We begin by noting that if \(U + \phi_{\varepsilon,\delta}\) 
    	is a radial solution to \eqref{elproblemaDelta2}, it must 
    	also satisfy
        \begin{equation}\label{eq:redu}
            \Pi(U + \phi_{\varepsilon,\delta} - 
            \mathcal{I}_\delta(g_{\varepsilon,\delta}(U + 
            \phi_{\varepsilon,\delta}))) = 0,
        \end{equation}
        where $g_{\varepsilon,\delta}$ was given in \eqref{la_g}. Then, our initial objective is to prove the existence of 
        a function \(\phi_{\varepsilon,\delta}\) such that 
        \(U + \phi_{\varepsilon,\delta}\) solves \eqref{eq:redu}. 
        To achieve this, we define the operator 
        \(T_{\varepsilon,\delta}\colon K \to K\) by
        \[
            T_{\varepsilon,\delta}(\phi) \coloneqq 
            \widetilde{L}^{-1}_\delta 
            \bigg(\Pi\Big(\mathcal{I}_\delta
            \big(f_\varepsilon(U + \phi) - f_0(U) - f_0^\prime(U) 
            \phi - \delta^{2(1-s)}(-\Delta)^s U\big)\Big)\bigg).
        \]
        Taking into account the decay of \eqref{bubble}, by using \cite[Lemma 2.1]{BonforteVa} with $\alpha=N-1$ we get that  $U\in\mathcal{D}^{1,2}(\mathbb{R}^{n})$. Thus, by \eqref{Condicion1} is clear that $(-\Delta)^sU\in \mathcal{X}$ so $T_{\varepsilon,\delta}$ is well defined.

        Observe that if \(\phi_{\varepsilon,\delta}\) is a 
        fixed point of \(T_{\varepsilon,\delta}\), 
        then \(U + \phi_{\varepsilon,\delta}\) is a solution of 
        equation \eqref{eq:redu}. Thus, our first task reduces 
        to showing that \(T_{\varepsilon,\delta}\) possesses a fixed 
        point.

        Once we establish the existence of a fixed point 
        \(\phi_{\varepsilon,\delta}\) for \(T_{\varepsilon,\delta}\), 
        our next objective will be to demonstrate that, for 
        appropriate values of \(\delta\) and \(\varepsilon\), 
        we have
        \begin{equation}\label{eq:contraepsi}
            \langle U + \phi_{\varepsilon,\delta} - 
            \mathcal{I}_\delta(g_{\varepsilon,\delta}(U + 
            \phi_{\varepsilon,\delta})), \psi \rangle = 0.
        \end{equation}
        
        Finally by \eqref{eq:redu} and \eqref{eq:contraepsi}, we 
        will conclude that \(U + \phi_{\varepsilon,\delta}\) 
        satisfies equation \eqref{elproblemaDelta2} for 
        appropriate values of \(\delta\) and \(\varepsilon.\)
        
        \medskip
To prove the existence of a fixed point, we first get the following

        \begin{lemma}\label{lemma:TgoesBtoB} 
       		Let  \(\alpha\in(0,1)\) and 
       		$\phi\in B^\alpha_{\varepsilon,\delta}:=\{{\phi\in {K}}:\, \|\phi\|\le (\varepsilon+\delta^{2(1-s)})^\alpha\}$. Then
       		\[
       			\|T_{\delta,\epsilon}(\phi)\|\le  
       			(\varepsilon+\delta^{2(1-s)})^{\alpha},
       		\]
       		for \(\varepsilon,\delta\)  small enough. That is, $T_{\delta,\epsilon}|_{B^\alpha_{\varepsilon,\delta}}
            \colon
            B^\alpha_{\varepsilon,\delta}
            \to B^\alpha_{\varepsilon,\delta}.
$
        \end{lemma}
        \begin{proof}
       		Let \(\alpha\in (0,1).\) Given \(\phi\in K\) , by Remark \ref{remark:Ideltacont} and Lemma \ref{lemaeq:ker} is clear that
       		\begin{align*}
       			\|T_{\delta,\epsilon}(\phi)\|&\le 
       				C\|\mathcal{I}_\delta(
            			f_\varepsilon(U+\phi)-f_0(U)-f_0^\prime(U)\phi-
            					\delta^{2(1-s)}(-\Delta)^s U))\|\\
            			&\le C\|\mathcal{I}_\delta(
           				 f_\varepsilon(U+\phi)-f_\varepsilon(U)
           				 -f_\varepsilon^\prime(U)\phi)\|
           				 + C\|\mathcal{I}_\delta(f_\varepsilon(U)-f_0(U))\|\\
          				  &+C\|\mathcal{I}_\delta(f_\varepsilon^\prime(U)\phi
          				  -f_0^\prime(U)\phi)\|
            				+\delta^{2(1-s)}C\|\mathcal{I}_\delta((-\Delta)^s U))\|\\
				&\le 
       					C\|\mathcal{I}_\delta(
           				 f_\varepsilon(U+\phi)-f_\varepsilon(U)
           				-f_\varepsilon^\prime(U)\phi\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}\\
           				&+ C\|f_\varepsilon(U)-f_0(U)\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}
          				  +C\|f_\varepsilon^\prime(U)\phi
          				  -f_0^\prime(U)\phi\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}\\
            				&+\delta^{2(1-s)}C
            		\|(-\Delta)^s U)\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}.\\
       		\end{align*}
       		Then, by H\''older inequality and {using \cite[Remark 2.1 and Remark 2.22]{MR1938656}}, we get
       		\[
       			\|T_{\delta,\epsilon}(\phi)\|\le 
       					C\left(\|\phi\|^{p_1-\varepsilon}
           				+\varepsilon+\varepsilon \|\phi\|
            				+\delta^{2(1-s)}\right),
       		\]
       		for \(\varepsilon\) small enough. Imposing now the condition  \(\|\phi\|\le (\varepsilon+\delta^{2(1-s)})^\alpha\) we conclude
       		\begin{align*}
					\|T_{\delta,\epsilon}(\phi)\|&\le 
       					C\left((\varepsilon+\delta^{2(1-s)})^{\alpha(p_1-\varepsilon)}
           				+(\varepsilon+\delta^{2(1-s)})^{\alpha+1}
            				+\varepsilon+\delta^{2(1-s)}\right)\\
            				&\le (\varepsilon+\delta^{2(1-s)})^{\alpha},
			\end{align*}
       		due to \(p_1-\varepsilon>1\) and the fact that \(\alpha\in(0,1)\).
        \end{proof}
       
        Our second step is to prove the next.    
        \begin{lemma}\label{lemma:Tiscontraction} 
            For \(\varepsilon\)
            small enough we have that 
            \(T_{\delta,\epsilon}|_{B^\alpha_{\varepsilon,\delta}}\)
            is a contraction mapping.
        \end{lemma}
        \begin{proof}
            Let \(\phi_1,\phi_2\in B^\alpha_{\varepsilon,\delta}.\)
            Then
            \begin{align*}
	            \|T_{\delta,\epsilon}(\phi_1)-T_{\delta,\epsilon}&(\phi_2)\|
	            \le C\|\mathcal{I}_\delta(f_\varepsilon(U+\phi_1)-
	                f_\varepsilon(U+\phi_2)-f_0^\prime(U)(\phi_1-\phi_2))\|\\
	            &\le C\|\mathcal{I}_\delta(f_\varepsilon(U+\phi_1)-
	                f_\varepsilon(U+\phi_2)-f_\varepsilon^\prime(U+\phi_2)
	                (\phi_1-\phi_2)))\|\\
	                &+C\|\mathcal{I}_\delta((f_\varepsilon^\prime(U+\phi_2)-
	                f_\varepsilon^\prime(U))(\phi_1-\phi_2))\|\\
	                &+C\|\mathcal{I}_\delta((f_\varepsilon^\prime(U)-f'_0(U))
	                (\phi_1-\phi_2))\|.
            \end{align*}
            Now by Remark \ref{remark:Ideltacont} and 
            \cite[Remark 2.1 and Remark 2.2.]{MR1938656}, we get
            \begin{align*}
	            &\|T_{\delta,\epsilon}(\phi_1)-T_{\delta,\epsilon}(\phi_2)\|\\
	            &\le C\left(\|\phi_1-\phi_2\|^{p_1-\varepsilon}+
	            \|\phi_2\|^{p_1-1-\varepsilon}\|\phi_1-\phi_2\|+{
	            \varepsilon\|\phi_1-\phi_2\|}
	            \right)\\
	            &\le D \|\phi_1-\phi_2\|,
            \end{align*}
            for some positive constant \(D<1\) due to \(p_1-\varepsilon>1.\)
        \end{proof}
        
        Then, by Lemmas \ref{lemma:TgoesBtoB} and \ref{lemma:Tiscontraction} we can conclude the existence of a fix point. That is.
         
        \begin{lemma}\label{eq:puntofijo}
            For \(\varepsilon\) and $\delta$ small enough there exists
           $\phi_{\varepsilon,\delta}\in {B^\alpha_{\varepsilon,\delta}}$
            such that 
          $$
                T_{\delta,\epsilon}(\phi_{\varepsilon,\delta})
            =\phi_{\varepsilon,\delta}.
            $$
        \end{lemma} 
        
        As we mentioned before, our final goal is to show that for 
        appropriate values of \(\delta\) and \(\varepsilon\), 
        we have \eqref{eq:contraepsi}. From now on, we take \(\varepsilon\) such that \(p_1-\varepsilon>1\), (i.e. $\varepsilon<\frac{4}{n-2}$), \(\alpha>\frac{1}{p_1-\varepsilon}\) and we consider $\phi_{\varepsilon,\delta}\in B^\alpha_{\varepsilon,\delta}$. 
        
        \begin{lemma} \label{lemaimp}
            For  $0< \varepsilon,\delta<1$ small enough 
            we get
            \begin{align*}
                \langle U + \phi_{\varepsilon,\delta} -
                \mathcal{I}_\delta(g_{\varepsilon,\delta}(U + 
                \phi_{\varepsilon,\delta})), \psi \rangle =
                \varepsilon\int_{\mathbb{R}^n}\log(U)U^{p_1}\psi \, dx +
                \delta^{2(1-s)}\langle U,\psi\rangle_s\\
                 +o(\varepsilon+\delta^{2(1-s)}).
            \end{align*}
        \end{lemma}
       
        \begin{proof}
            Let
            \[
                W_{\varepsilon,\delta}\coloneqq 
                \mathcal{I}_\delta(g_{\varepsilon,\delta}(U + 
                \phi_{\varepsilon,\delta})).
            \]
            Then, using the equations satisfy by \(U\) and \(\psi,\) we get
            \begin{align*}
	            &\langle U + \phi_{\varepsilon,\delta} - 
                \mathcal{I}_\delta(g_{\varepsilon,\delta}(U + 
                \phi_{\varepsilon,\delta})), \psi \rangle = 
                \langle U + \phi_{\varepsilon,\delta} - 
                W_{\varepsilon,\delta}, \psi \rangle\\
                &= \langle U , \psi \rangle+
                 \langle  \phi_{\varepsilon,\delta} , \psi \rangle
                 - \langle W_{\varepsilon,\delta}, \psi \rangle\\
                 &=\int_{\mathbb{R}^n}\left(f_0(U)+f^\prime_0(U) 
                 \phi_{\varepsilon,\delta}-
                 f_\varepsilon(U+\phi_{\varepsilon,\delta})  \right)\psi \, dx\\
                 &\qquad\qquad
                 +\delta^{2(1-s)}\int_{\mathbb{R}^n}(W_{\varepsilon,\delta}-U)\psi \, dx 
                 +\delta^{2(1-s)}\langle W_{\varepsilon,\delta},\psi\rangle_s
                 -\delta^{2(1-s)}\int_{\mathbb{R}^n}\phi_{\varepsilon,\delta}\psi \, dx\\
                 &=\int_{\mathbb{R}^n}\left(f_0(U)+f^\prime_0(U) 
                 \phi_{\varepsilon,\delta}-
                 f_\varepsilon(U+\phi_{\varepsilon,\delta})  \right)\psi \, dx
                 +\delta^{2(1-s)}\int_{\mathbb{R}^n}(W_{\varepsilon,\delta}-U)\psi \, dx\\
                 &\qquad\qquad
                 +\delta^{2(1-s)}\langle W_{\varepsilon,\delta}-U,\psi\rangle_s
                 +\delta^{2(1-s)}\langle U,\psi\rangle_s
                 -\delta^{2(1-s)}\int_{\mathbb{R}^n}\phi_{\varepsilon,\delta}\psi \, dx.
            \end{align*}
            Observe that
             \begin{align*}
	            \int_{\mathbb{R}^n}\big(f_0(U)+f^\prime_0(U)& 
                 \phi_{\varepsilon,\delta}-
                 f_\varepsilon(U+\phi_{\varepsilon,\delta})  \big)\psi \, dx=\\
                 =& \int_{\mathbb{R}^n}\left(f_\varepsilon(U)
                 +f^\prime_\varepsilon(U) 
                 \phi_{\varepsilon,\delta}-
                 f_\varepsilon(U+\phi_{\varepsilon,\delta})  \right)\psi \, dx\\
                 +&\int_{\mathbb{R}^n}\left(f_0(U)-f_\varepsilon(U)\right)
                 \psi \, dx+
                 \int_{\mathbb{R}^n}\left(f_0^\prime(U)
                 -f_\varepsilon^\prime(U)\right)
                 \psi \, dx.
            \end{align*}
            Following the computations done in \cite[(3.20)--(3.23)]{MR1938656}, using that \(\alpha>\frac1{p_1-\varepsilon}\) 
            and $ {(p_1-\varepsilon)\frac{2n}{n+2}}\in \left(\frac{2n}{n+2},\frac{2n}{n-2}\right)$
            for $\varepsilon$ sufficiently small, we get
            \[
                \int_{\mathbb{R}^n}\big(f_0(U)+f^\prime_0(U) 
                 \phi_{\varepsilon,\delta}-
                 f_\varepsilon(U+\phi_{\varepsilon,\delta})  \big)\psi \, dx
                 =\varepsilon\int_{\mathbb{R}^n}\log(U)U^{p_1}\psi \, dx +
                 o(\varepsilon+\delta^{2(1-s)}).
            \]
           
            On the other hand, since \(p_1-\varepsilon>1\), applying Lemma \ref{lemma:ordenconv}
            and Remark \ref{remark.psi},  we get
            \begin{align*}
	            &\Bigg|\delta^{2(1-s)}\int_{\mathbb{R}^n}(w_{\varepsilon,\delta}-U)\psi \, dx
                 +\delta^{2(1-s)}\langle w_{\varepsilon,\delta}-U,\psi\rangle_s
                 -\delta^{2(1-s)}\int_{\mathbb{R}^n}\phi_{\varepsilon,\delta}\psi \, dx\Bigg|\\
                 &\le 
                 C\left(\delta^{2(1-s)}\|w_{\varepsilon,\delta}-U\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}
                 +\delta^{2(1-s)}\|w_{\varepsilon,\delta}-U\|_{s,2}
                 +\delta^{2(1-s)}\|\phi_{\varepsilon,\delta}\|_{L^{\frac{2n}{n+2}}(\mathbb{R}^n)}\right)\\
                 &\le C\delta^{2(1-s)}
                 \left(\|\phi_{\varepsilon,\delta}\|^{p_1-\varepsilon}+
                  \|\phi_{\varepsilon,\delta}\|
				+\varepsilon+\delta^{2(1-s)}\right)\\
				&\le 
				C\delta^{2(1-s)}\left(\varepsilon +\delta^{2(1-s)}\right)^{\alpha}.
            \end{align*}
        \end{proof}
        
        We can finally prove our main result.
        \begin{proof}[Proof of Theorem \ref{ElTeorema}]
        	Let be $\varepsilon>0$ and $\lambda>0$. Applying Lemma \ref{lemaimp} with $\delta^{2(1-s)}:=\varepsilon\lambda$, we have that
        	\[
        		\langle U + \phi_{\varepsilon,\delta} -
        		\mathcal{I}_\delta(g_{\varepsilon,\delta}(U + 
        		\phi_{\varepsilon,\delta})), \psi \rangle =
        		\varepsilon\left(A+\lambda B
        		+\frac{o(\varepsilon+\varepsilon\lambda)}{\varepsilon}\right)=:\varepsilon F(\varepsilon,\lambda).
        	\]
       Since by Lemma \ref{taqui}, 
        	$$
        		A\coloneqq\int_{\mathbb{R}^n}\log(U)U^{p_1}\psi \, dx >0,\text{ and } B\coloneqq\langle U,\psi\rangle_s<0,
        	$$
        	we can take $\lambda_0:=-A/B>0$, that implies
	$$F(\varepsilon,\lambda_0)=(1+\lambda_0)\frac{o(\varepsilon(1+\lambda_0))}{\varepsilon(1+\lambda_0)}=o(1)\, \mbox{ (with respect to $\varepsilon$)}.$$
	That is, \(F(\varepsilon,\lambda_0)=0\) for 
        	$\varepsilon$ small enough, as wanted.
        \end{proof}
        
        \section*{Acknowledgments}
         All the authors were partially supported by the project {\it An\'alisis de Fourier y Ecuaciones no locales en Derivadas Parciales} Grant PID2023-148028NB-I00 founded by MCIN/ AEI/10.13039/501100011033/ FEDER, UE.
         AQ was also partially supported by FONDECYT Grant 1231585.
          LMDP was also supported by ANII  under the programs  “Movilizaciones AMSUD 2023-01” 
         (MOV CO 2023 1 1012406 and MOV CO 9 101336) and ANII FCE 3 2024 1 181302.
 
 \bibliographystyle{abbrv}
\bibliography{Bibliografia}

\end{document}